x-5(x+2)=x+3(3-2x)
Answer by checkley75(3666)

X-5(X+2)=X+3(3-2X)
X-5X-10=X+9-6X
X-5X-X+6X=9+10
X=19 ANSWER.
PROOF
19-5(19+2)=19+3(3-2*19)
19-5(21)=19+3(3-38)
19-105=19+3*-35
86=19-105
86=86